# Free Descriptions of Convex Sets

### v2

@@ Abstract
Convex sets arising in a variety of applications are well-defined for every relevant dimension. Examples include the simplex and the spectraplex that correspond to probability distributions and to quantum states; combinatorial polytopes and their relaxations such as the cut polytope and the elliptope in integer programming; and unit balls of regularizers such as the $\ell_{p}$ and Schatten norms in inverse problems. Moreover, these sets are often specified using conic descriptions that can be obviously instantiated in any dimension. We develop a systematic framework to study such free descriptions of convex sets. We show that free descriptions arise from a recently-identified phenomenon in algebraic topology called representation stability, which relates invariants across dimensions in a sequence of group representations. Our framework yields structural results for free descriptions pertaining to the relations between the sets they describe across dimensions, extendability of a single set in a given dimension to a freely-described sequence, and continuous limits of such sequences. We also develop a procedure to obtain parametric families of freely-described convex sets whose structure is adapted to a given application; illustrations are provided via examples that arise in the literature as well as new families that are derived using our procedure. We demonstrate the utility of our framework in two contexts. First, we develop an algorithm for a free analog of the convex regression problem, where a convex function is fit to input-output data; by searching over our parametric families, we can fit a function to low-dimensional inputs and extend it to any other dimension. Second, we prove that many sequences of symmetric conic programs can be solved in constant time, which unifies and strengthens several results in the literature.
@@ 1 Introduction
Convex sets play a central role in numerous areas of the mathematical sciences such as optimization, statistical inference, control, inverse problems, and information theory. The convex sets arising in all these domains are often well-defined in every relevant dimension. Indeed, unit balls of standard regularizers used in inverse problems (e.g., $\ell_{p}$ or Schatten norms) are defined for vectors and matrices of any size. Polytopes associated to graph problems and their relaxations (e.g., the cut polytope and the elliptope approximating it) are defined for graphs of any size. Information-theoretic quantities (e.g., relative entropy) as well as their quantum analogues, are defined for distributions and states on any number of (qu)bits. Consequently, all these convex sets and many others should be viewed not as single sets but as sequences indexed by dimension. Furthermore, there is often a single “free” description of all the sets in such a sequence, as the following example illustrates.
@@
[…]
@@
Note that all of the above descriptions are clearly defined for any relevant dimension. Indeed, these descriptions are composed of elements such as $\mathbbm{1}_{n}$ and $I_{n}$, linear maps such as $\mathrm{diag}(\cdot)$ and $\mathrm{Tr}(\cdot)$, and inequalities such as $\geq$ and $\succeq$, all of which are “free” in the sense that they are well-defined in every relevant dimension. In this paper, we develop a systematic framework to study such freely-described convex sets.
@@
    The sequence $\mathcal{C}_{n}=\Bigg{\{}X\in\mathbb{S}^{n}:\frac{\mathbbm{1}_{n}^{\top}X\mathbbm{1}_{n}}{n^{2}}L_{1}\otimes\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}+\frac{\mathrm{Tr}(X)}{n}L_{2}\otimes\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}+L_{3}\otimes\frac{1}{n}\left(X\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}+\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}X\right)+L_{4}\otimes\left(\mathrm{diag}(X)\mathbbm{1}_{n}^{\top}+\mathbbm{1}_{n}\mathrm{diag}(X)^{\top}\right)+L_{5}\otimes X+L_{6}\otimes\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}+L_{7}\otimes(nI_{n})\succeq 0\Bigg{\}}$ is parametrized by $L_{1},\ldots,L_{7}\in\mathbb{S}^{k}$. We obtain this sequence from our framework in Section 4.7 by viewing a matrix $X\in\mathbb{S}^{n}$ representing a weighted graph as a step graphon [15], and deriving a parametric family of free descriptions extending continuously to more general spaces of graphons.
@@
Note that the parametric free descriptions in Example 1.2 are composed of linear combinations of the free elements and linear maps that appear in the descriptions of Example 1.1. In fact, the parameters above are simply the coefficients in these linear combinations. Thus, we can obtain finitely-parametrized families of freely-described convex sets by deriving finite-dimensional spaces of free elements.
The central thesis of this paper is that the free elements which constitute free descriptions arise from a recently-identified phenomenon in algebraic topology called representation stability, which relates invariants across dimensions in a sequence of group representations [16]. We use these relations to formally define the free elements that constitute our free descriptions, and investigate various structural properties of such descriptions as well as the associated sequences of optimization problems. Our definitions together with results from representation stability also yield finite-dimensional spaces of free elements, which we use to derive parametric families of freely-described convex sets adapted to specific applications and to fit elements of these families to data. We outline the contributions of this paper in the remainder of the introduction. We do not assume prior familiarity with representation stability, and the relevant background is presented in Section 2.
@@ 1.1 Our Framework and Contributions
This paper consists of four main contributions. First, we formally define free descriptions of convex sets by generalizing the insights derived from Examples 1.1 and 1.2 using representation stability. Second, we prove structural results pertaining to freely-described sequences of convex sets, the relationship between sets in such a sequence in different dimensions, and limits of such sequences. Our proofs combine concepts from convex analysis and representation theory. Third, we derive consequences of our framework for invariant conic programs, showing that sequences of such programs of growing dimension can often be solved in constant time. Our results unify and generalize existing work in the literature. Finally, we apply our framework and its structural results to derive an algorithm for fitting a freely-described convex function to given data in different dimensions, a problem we call free convex regression. The latter two contributions are aimed at addressing our mathematical and algorithmic motivations above.
@@ 1.1.1 Freely-described Convex Sets
To formalize free descriptions of convex sets, we begin by briefly reviewing conic descriptions, which express a convex set as an affine section of a convex cone. Conic descriptions are the most popular approach to specifying convex sets in the optimization literature, and they have played a central role in the development of modern convex optimization [17]. Indeed, we often classify convex sets based on their conic descriptions—polyhedra are affine sections of nonnegative orthants and spectrahedra are affine sections of PSD cones. Formally, if $\mathbb{V},\mathbb{W},\mathbb{U}$ are (finite-dimensional) vector spaces and $\mathcal{K}\subseteq\mathbb{U}$ is a convex cone, then a convex subset $\mathcal{C}\subseteq\mathbb{V}$ can be described using linear maps $A\colon\mathbb{V}\to\mathbb{U}$, $B\colon\mathbb{W}\to\mathbb{U}$ and a vector $u\in\mathbb{U}$ as follows:
@@
Each sequence of convex sets $\mathcal{C}_{n}\subseteq\mathbb{V}_{n}$ in Examples 1.1 and 1.2 is given by (Conic) for suitable sequences of description spaces $\mathbb{W}_{n},\mathbb{U}_{n}$, of vectors $u_{n}\in\mathbb{U}_{n}$, of linear maps $A_{n}\colon\mathbb{V}_{n}\rightarrow\mathbb{U}_{n}$, $B_{n}\colon\mathbb{W}_{n}\rightarrow\mathbb{U}_{n}$, and of convex cones $\mathcal{K}_{n}\subseteq\mathbb{U}_{n}$. In particular, all the sequences of cones $\mathcal{K}_{n}$ in those examples are standard sequences such as the nonnegative orthants and PSD cones. Expressing convex sets in terms of such standard sequences has the benefit that optimization over these sets can be performed using standard off-the-shelf software. Thus, all the sequences of cones, as well as the vector spaces containing them, appearing in this paper come from such standard sequences.
Having fixed standard sequences of vector spaces and cones, we seek to formalize the free vectors $u_{n}$ and linear maps $A_{n},B_{n}$ that are defined for any dimension, such as those appearing in Examples 1.1 and 1.2. Furthermore, we wish to obtain finite-dimensional spaces of these free elements, which yield finitely-parametrized families of free descriptions. To that end, we observe that these free elements are sequences of invariants under sequences of groups related in a particular way across dimensions. For example, the all-1’s vector $\mathbbm{1}_{n}$ of length $n$ is invariant under the group of permutations on $n$ letters acting by permuting coordinates. Further, the all-1’s vectors of different lengths are related to each other: extracting the first $n$ entries of $\mathbbm{1}_{n+1}$ yields $\mathbbm{1}_{n}$. Similarly, the $n\times n$ identity matrix $I_{n}$ is invariant under the orthogonal group of size $n$ acting by conjugation, and extracting the top left $n\times n$ submatrix of $I_{n+1}$ yields $I_{n}$. Thus, to give a formal definition for free vectors and linear maps, we consider sequences of groups acting on sequences of vector spaces, and we require the spaces in the sequence to be related to each other – specifically, we embed lower-dimensional spaces into higher-dimensional ones and project higher-dimensional spaces onto lower-dimensional ones. Such sequences of group representations are called consistent sequences and they were first defined in the seminal paper [16] introducing representation stability.
@@
As $\varphi_{n}$ is an isometry, its adjoint $\varphi_{n}^{*}$ defines the orthogonal projection $\mathcal{P}_{\mathbb{V}_{n}}$ of higher-dimensional spaces onto lower-dimensional ones. We now formally define free vectors or linear maps as sequences of invariants projecting onto each other.
@@
Having formalized freely-described vectors and linear maps, we are ready to define free descriptions of convex sets. These are just sequences of conic descriptions specified using freely-described elements.
Definition 1.6 (Free conic descriptions).
@@
is called free if $\{A_{n}\},\{B_{n}\}$, and $\{u_{n}\}$ are freely-described elements of the consistent sequences $\mathscr{V}\otimes\mathscr{U}$, $\mathscr{W}\otimes\mathscr{U}$, and $\mathscr{U}$, respectively.
All the descriptions in Examples 1.1 and 1.2 become free when the relevant sequences of vector spaces are endowed with natural consistent sequence structure. Moreover, when $\mathscr{V}\otimes\mathscr{U},\mathscr{W}\otimes\mathscr{U},\mathscr{U}$ are all finitely-generated, Definition 1.6 yields a finitely-parametrized family of free descriptions, obtained by choosing bases for the spaces of freely-described $\{A_{n}\},\{B_{n}\},\{u_{n}\}$ and viewing the coefficients in these bases as the parameters. These parametric families generalize Example 1.2. More broadly, they can be adapted to specific applications via the choice of embeddings and group actions involved in the consistent sequences, which formally relate instances of different sizes and their symmetries.
We can use free descriptions to describe convex functions as well as sets, using the many correspondences between convex sets and functions from convex analysis. For example, given a freely-described sequence of convex sets we can consider the sequence of their support and gauge functions, and given a sequence of convex functions we can consider free descriptions for the sequence of their epigraphs [18, §4]. Thus, Definition 1.6 yields finitely-parametrized infinite sequences of convex functions as well.
@@
$\mathcal{C}_{n}$ =$=\Bigg{\{}X\in\mathbb{S}^{n}:\alpha_{1}(\mathbbm{1}_{n}^{\top}X\mathbbm{1}_{n})\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}+\alpha_{2}(\mathbbm{1}_{n}^{\top}X\mathbbm{1}_{n})I_{n}+\alpha_{3}\mathrm{Tr}(X)\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}+\alpha_{4}\mathrm{Tr}(X)I_{n}+\alpha_{5}\Big{(}X\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}+\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}X\Big{)}$ (1)
@@
1.1.2 Structural Aspects of Free Descriptions
@@
  1. missingfnum@enumiitemIntersection compatibility
@@
  2. missingfnum@enumiitemProjection compatibility
@@
Our first structural result gives conditions under which free descriptions certify compatibility, i.e., under which the sequence of sets derived from these descriptions is evidently compatible. Consider a freely-described sequence $\{\mathcal{C}_{n}\}$ of convex sets (ConicSeq). Assuming that the sequence of cones $\{\mathcal{K}_{n}\}$ underlying the description of $\{\mathcal{C}_{n}\}$ is both intersection and projection compatible, we prove in Proposition 3.2 that the sequence of convex sets $\{\mathcal{C}_{n}\}$ satisfies compatibility conditions if the underlying freely-described elements in (ConicSeq) lie in a convex cone. This result serves as the foundation for our subsequent developments—we use it to show in Section 4 that many sets arising in applications naturally admit free descriptions certifying their compatibility. We also use this result in Section 6 to design an algorithm fitting a freely-described and compatible sequence of sets to data.
Our second structural result is central to the development of our computational framework in Section 6 and it addresses the following question. Given a convex set with a conic description (Conic) in a fixed dimension $n_{0}$, when does this description extend to a free description of a sequence of sets satisfying compatibility conditions? As a concrete example, suppose we are given a convex relaxation for a combinatorial optimization problem, or a convex regularizer for an inverse problem, in a particular problem dimension; under what conditions can these be extended to problems in any desired dimension? Leveraging our preceding result along with properties of presentation degrees of consistent sequences, we give conditions in Theorem 3.5 on the dimension $n_{0}$ and the conic description in that dimension ensuring the desired extendability. We use this result in Section 6 to computationally extend a set fitted to data in a fixed dimension to any other dimension.
@@
Consider the freely-described sequence in Example 1.2(c), which is intersection- and projection-compatible with respect to the graphon consistent sequence (see Section 4.7). The elements of $\mathbb{V}_{\infty}$ are step functions on $[0,1]^{2}$ known as step graphons, and we endow them with the $L_{\infty}$ norm as is common in the literature [22]. The completion $\overline{\mathbb{V}_{\infty}}$ then consists of certain piecewise-continuous graphons, and Theorem 3.6 shows that the limit of the sequence $\{\mathcal{C}_{n}\}$ in Example 1.2(c) with $L_{7}=0$ is
@@
Here positive-semidefiniteness of a matrix-valued function on $[0,1]^{2}$ is meant in the sense of matrix-valued kernels [23], see (21) for a definition. We require $L_{7}=0$ because the corresponding term does not extend continuously with respect to the $L_{\infty}$ norm. We derive this limiting description in Proposition 4.7.
We also apply Theorem 3.6 to permutahedra and Schur-Horn orbitopes in Section 4.3 to obtain descriptions of limits in analogy to Example 1.9. The resulting infinite-dimensional conic descriptions yield an infinite-dimensional generalization of the Schur-Horn theorem, which we give in Proposition 4.3. In fact, we obtain limiting descriptions and a Schur-Horn theorem more generally in approximately finite-dimensional (AF) algebras in Appendix C.

### v3

# Dimension-Free Descriptions of Convex Sets

@@ Abstract
Convex sets arising in a variety of applications are well-defined for every relevant dimension. Examples include the simplex and the spectraplex that correspond to probability distributions and to quantum states; combinatorial polytopes and their relaxations such as the cut polytope and the elliptope in integer programming; and unit balls of regularizers such as the $\ell_{p}$ and Schatten norms in inverse problems. Moreover, these sets are often specified using conic descriptions that can be obviously instantiated in any dimension. We develop a systematic framework to study such dimension-free descriptions of convex sets. We show that dimension-free descriptions arise from a recently-identified phenomenon in algebraic topology called representation stability, which relates invariants across dimensions in a sequence of group representations. Our framework yields structural results for dimension-free descriptions pertaining to the relations between the sets they describe across dimensions, extendability of a single set in a given dimension to a freely-described sequence, and continuous limits of such sequences. We also develop a procedure to obtain parametric families of freely-described convex sets whose structure is adapted to a given application; illustrations are provided via examples that arise in the literature as well as new families that are derived using our procedure. We demonstrate the utility of our framework in two contexts. First, we develop an algorithm for a dimension-free analog of the convex regression problem, where a convex function is fit to input-output data; by searching over our parametric families, we can fit a function to low-dimensional inputs and extend it to any other dimension. Second, we prove that many sequences of symmetric conic programs can be solved in constant time, which unifies and strengthens several results in the literature.
@@ 1 Introduction
Convex sets play a central role in numerous areas of the mathematical sciences such as optimization, statistical inference, control, inverse problems, and information theory. The convex sets arising in all these domains are often well-defined in every relevant dimension. Indeed, unit balls of standard regularizers used in inverse problems (e.g., $\ell_{p}$ or Schatten norms) are defined for vectors and matrices of any size. Polytopes associated to graph problems and their relaxations (e.g., the cut polytope and the elliptope approximating it) are defined for graphs of any size. Information-theoretic quantities (e.g., relative entropy) as well as their quantum analogues, are defined for distributions and states on any number of (qu)bits. Consequently, all these convex sets and many others should be viewed not as single sets but as sequences indexed by dimension. Furthermore, there is often a single “dimension-free” description of all the sets in such a sequence, as the following example illustrates.
@@
    The simplex in $n$ dimensions is $\Delta^{n-1}=\{x\in\mathbb{R}^{n}:x\geq 0,\ \mathbbm{1}_{n}^{\top}x=1\}$, which is the set of probability distributions over $n$ items. Here $x\geq 0$ denotes an entrywise nonnegative vector $x$, and $\mathbbm{1}_{n}\in\mathbb{R}^{n}$ is the vector of all-1’s.
@@
Note that all of the above descriptions are clearly defined for any relevant dimension. Indeed, these descriptions are composed of elements such as $\mathbbm{1}_{n}$ and $I_{n}$, linear maps such as $\mathrm{diag}(\cdot)$ and $\mathrm{Tr}(\cdot)$, and inequalities such as $\geq$ and $\succeq$, all of which are “dimension-free” in the sense that they are well-defined in every relevant dimension. In this paper, we develop a systematic framework to study such freely-described convex sets.
@@
    The sequence $\mathcal{C}_{n}=\Bigg\{X\in\mathbb{S}^{n}:\frac{\mathbbm{1}_{n}^{\top}X\mathbbm{1}_{n}}{n^{2}}L_{1}\otimes\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}+\frac{\mathrm{Tr}(X)}{n}L_{2}\otimes\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}+L_{3}\otimes\frac{1}{n}\left(X\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}+\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}X\right)+L_{4}\otimes\left(\mathrm{diag}(X)\mathbbm{1}_{n}^{\top}+\mathbbm{1}_{n}\mathrm{diag}(X)^{\top}\right)+L_{5}\otimes X+L_{6}\otimes\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}+L_{7}\otimes(nI_{n})\succeq 0\Bigg\}$ is parametrized by $L_{1},\ldots,L_{7}\in\mathbb{S}^{k}$. We obtain this sequence from our framework in Section 4.7 by viewing a matrix $X\in\mathbb{S}^{n}$ representing a weighted graph as a step graphon [15], and deriving a parametric family of dimension-free descriptions extending continuously to more general spaces of graphons.
@@
Note that the parametric dimension-free descriptions in Example 1.2 are composed of linear combinations of the dimension-free elements and linear maps that appear in the descriptions of Example 1.1. In fact, the parameters above are simply the coefficients in these linear combinations. Thus, we can obtain finitely-parametrized families of freely-described convex sets by deriving finite-dimensional spaces of dimension-free elements.
The central thesis of this paper is that the dimension-free elements which constitute dimension-free descriptions arise from a recently-identified phenomenon in algebraic topology called representation stability, which relates invariants across dimensions in a sequence of group representations [16]. We use these relations to formally define the dimension-free elements that constitute our dimension-free descriptions, and investigate various structural properties of such descriptions as well as the associated sequences of optimization problems. Our definitions together with results from representation stability also yield finite-dimensional spaces of dimension-free elements, which we use to derive parametric families of freely-described convex sets adapted to specific applications and to fit elements of these families to data. We outline the contributions of this paper in the remainder of the introduction. We do not assume prior familiarity with representation stability, and the relevant background is presented in Section 2.
@@ 1.1 Our Framework and Contributions
This paper consists of four main contributions. First, we formally define dimension-free descriptions of convex sets by generalizing the insights derived from Examples 1.1 and 1.2 using representation stability. Second, we prove structural results pertaining to freely-described sequences of convex sets, the relationship between sets in such a sequence in different dimensions, and limits of such sequences. Our proofs combine concepts from convex analysis and representation theory. Third, we derive consequences of our framework for invariant conic programs, showing that sequences of such programs of growing dimension can often be solved in constant time. Our results unify and generalize existing work in the literature. Finally, we apply our framework and its structural results to derive an algorithm for fitting a freely-described convex function to given data in different dimensions, a problem we call dimension-free convex regression. The latter two contributions are aimed at addressing our mathematical and algorithmic motivations above.
@@ 1.1.1 Freely-described Convex Sets
To formalize dimension-free descriptions of convex sets, we begin by briefly reviewing conic descriptions, which express a convex set as an affine section of a convex cone. Conic descriptions are the most popular approach to specifying convex sets in the optimization literature, and they have played a central role in the development of modern convex optimization [17]. Indeed, we often classify convex sets based on their conic descriptions—polyhedra are affine sections of nonnegative orthants and spectrahedra are affine sections of PSD cones. Formally, if $\mathbb{V},\mathbb{W},\mathbb{U}$ are (finite-dimensional) vector spaces and $\mathcal{K}\subseteq\mathbb{U}$ is a convex cone, then a convex subset $\mathcal{C}\subseteq\mathbb{V}$ can be described using linear maps $A\colon\mathbb{V}\to\mathbb{U}$, $B\colon\mathbb{W}\to\mathbb{U}$ and a vector $u\in\mathbb{U}$ as follows:
@@
Each sequence of convex sets $\mathcal{C}_{n}\subseteq\mathbb{V}_{n}$ in Examples 1.1 and 1.2 is given by (Conic) for suitable sequences of description spaces $\mathbb{W}_{n},\mathbb{U}_{n}$, of vectors $u_{n}\in\mathbb{U}_{n}$, of linear maps $A_{n}\colon\mathbb{V}_{n}\rightarrow\mathbb{U}_{n}$, $B_{n}\colon\mathbb{W}_{n}\rightarrow\mathbb{U}_{n}$, and of convex cones $\mathcal{K}_{n}\subseteq\mathbb{U}_{n}$. The vector spaces and cones are typically fixed in advance and are dictated by the application domain. Having fixed the vector spaces and cones, we seek to formalize the vectors $u_{n}$ and linear maps $A_{n},B_{n}$ that are defined for any dimension, such as those appearing in Examples 1.1 and 1.2. Furthermore, we wish to obtain finite-dimensional spaces of these dimension-free elements, which yield finitely-parametrized families of dimension-free descriptions. To that end, we observe that these dimension-free elements are sequences of invariants under sequences of groups related in a particular way across dimensions. For example, the all-1’s vector $\mathbbm{1}_{n}$ of length $n$ is invariant under the group of permutations on $n$ letters acting by permuting coordinates. Further, the all-1’s vectors of different lengths are related to each other: extracting the first $n$ entries of $\mathbbm{1}_{n+1}$ yields $\mathbbm{1}_{n}$. Similarly, the $n\times n$ identity matrix $I_{n}$ is invariant under the orthogonal group of size $n$ acting by conjugation, and extracting the top left $n\times n$ submatrix of $I_{n+1}$ yields $I_{n}$. Thus, to give a formal definition for dimension-free vectors and linear maps, we consider sequences of groups acting on sequences of vector spaces, and we require the spaces in the sequence to be related to each other – specifically, we embed lower-dimensional spaces into higher-dimensional ones and project higher-dimensional spaces onto lower-dimensional ones. Such sequences of group representations are called consistent sequences and they were first defined in the seminal paper [16] introducing representation stability.
@@
As $\varphi_{n}$ is an isometry, its adjoint $\varphi_{n}^{*}$ defines the orthogonal projection $\mathcal{P}_{\mathbb{V}_{n}}$ of higher-dimensional spaces onto lower-dimensional ones. We now formally define dimension-free vectors or linear maps as sequences of invariants projecting onto each other.
@@
Having formalized freely-described vectors and linear maps, we are ready to define dimension-free descriptions of convex sets. These are just sequences of conic descriptions specified using freely-described elements.
Definition 1.6 (Dimension-free conic descriptions).
@@
is called dimension-free if $\{A_{n}\},\{B_{n}\}$, and $\{u_{n}\}$ are freely-described elements of the consistent sequences $\mathscr{V}\otimes\mathscr{U}$, $\mathscr{W}\otimes\mathscr{U}$, and $\mathscr{U}$, respectively.
All the descriptions in Examples 1.1 and 1.2 become dimension-free when the relevant sequences of vector spaces are endowed with natural consistent sequence structure. Moreover, when $\mathscr{V}\otimes\mathscr{U},\mathscr{W}\otimes\mathscr{U},\mathscr{U}$ are all finitely-generated, Definition 1.6 yields a finitely-parametrized family of dimension-free descriptions, obtained by choosing bases for the spaces of freely-described $\{A_{n}\},\{B_{n}\},\{u_{n}\}$ and viewing the coefficients in these bases as the parameters. These parametric families generalize Example 1.2. More broadly, they can be adapted to specific applications via the choice of embeddings and group actions involved in the consistent sequences, which formally relate instances of different sizes and their symmetries.
We can use dimension-free descriptions to describe convex functions as well as sets, using the many correspondences between convex sets and functions from convex analysis. For example, given a freely-described sequence of convex sets we can consider the sequence of their support and gauge functions, and given a sequence of convex functions we can consider dimension-free descriptions for the sequence of their epigraphs [18, §4]. Thus, Definition 1.6 yields finitely-parametrized infinite sequences of convex functions as well.
@@
$\mathcal{C}_{n}$ $=\Bigg\{X\in\mathbb{S}^{n}:\alpha_{1}(\mathbbm{1}_{n}^{\top}X\mathbbm{1}_{n})\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}+\alpha_{2}(\mathbbm{1}_{n}^{\top}X\mathbbm{1}_{n})I_{n}+\alpha_{3}\mathrm{Tr}(X)\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}+\alpha_{4}\mathrm{Tr}(X)I_{n}+\alpha_{5}\Big(X\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}+\mathbbm{1}_{n}\mathbbm{1}_{n}^{\top}X\Big)$ (1)
@@
1.1.2 Structural Aspects of dimension-free descriptions
@@
  1. fnum@enumiitemIntersection compatibility
@@
  2. fnum@enumiitemProjection compatibility
@@
Our first structural result gives conditions under which dimension-free descriptions certify compatibility, i.e., under which the sequence of sets derived from these descriptions is evidently compatible. Consider a freely-described sequence $\{\mathcal{C}_{n}\}$ of convex sets (ConicSeq). Assuming that the sequence of cones $\{\mathcal{K}_{n}\}$ underlying the description of $\{\mathcal{C}_{n}\}$ is both intersection and projection compatible, we prove in Proposition 3.2 that the sequence of convex sets $\{\mathcal{C}_{n}\}$ satisfies compatibility conditions if the underlying freely-described elements in (ConicSeq) lie in a convex cone. This result serves as the foundation for our subsequent developments—we use it to show in Section 4 that many sets arising in applications naturally admit dimension-free descriptions certifying their compatibility. We also use this result in Section 6 to design an algorithm fitting a freely-described and compatible sequence of sets to data.
Our second structural result is central to the development of our computational framework in Section 6 and it addresses the following question. Given a convex set with a conic description (Conic) in a fixed dimension $n_{0}$, when does this description extend to a dimension-free description of a sequence of sets satisfying compatibility conditions? As a concrete example, suppose we are given a convex relaxation for a combinatorial optimization problem, or a convex regularizer for an inverse problem, in a particular problem dimension; under what conditions can these be extended to problems in any desired dimension? Leveraging our preceding result along with properties of presentation degrees of consistent sequences, we give conditions in Theorem 3.5 on the dimension $n_{0}$ and the conic description in that dimension ensuring the desired extendability. We use this result in Section 6 to computationally extend a set fitted to data in a fixed dimension to any other dimension.
@@
Consider the freely-described sequence in Example 1.2(c), which is intersection- and projection-compatible with respect to the graphon consistent sequence (see Section 4.7). The elements of $\mathbb{V}_{\infty}$ are step functions on $0,1]^{2}$ known as step graphons, and we endow them with the $L_{\infty}$ norm as is common in the literature [22]. The completion $\overline{\mathbb{V}_{\infty}}$ then consists of certain piecewise-continuous graphons, and Theorem 3.6 shows that the limit of the sequence $\{\mathcal{C}_{n}\}$ in Example 1.2(c) with $L_{7}=0$ is
@@
Here positive-semidefiniteness of a matrix-valued function on $0,1]^{2}$ is meant in the sense of matrix-valued kernels [23], see (21) for a definition. We require $L_{7}=0$ because the corresponding term does not extend continuously with respect to the $L_{\infty}$ norm. We derive this limiting description in Proposition 4.7.
We also apply Theorem 3.6 to permutahedra and Schur–Horn orbitopes in Section 4.3 to obtain descriptions of limits in analogy to Example 1.9. The resulting infinite-dimensional conic descriptions yield an infinite-dimensional generalization of the Schur–Horn theorem, which we give in Proposition 4.3. In fact, we obtain limiting descriptions and a Schur–Horn theorem more generally in approximately finite-dimensional (AF) algebras in Appendix C.